# Josephson diodes induced by the loop current states

### v1

@@ Acknowledgements.
- Hu et al. [2023] J.-X. Hu, Z.-T. Sun, Y.-M. Xie, and K. T. Law, Josephson diode effect induced by valley polarization in twisted bilayer graphene, Phys. Rev. Lett. 130, 266003 (2023).
@@
- Lu et al. [2023] B. Lu, S. Ikegaya, P. Burset, Y. Tanaka, and N. Nagaosa, Tunable josephson diode effect on the surface of topological insulators, Phys. Rev. Lett. 131, 096001 (2023).
@@
- Alvarado et al. [2023] M. Alvarado, P. Burset, and A. L. Yeyati, Intrinsic nonmagnetic josephson junctions in twisted bilayer graphene, Phys. Rev. Res. 5, L032033 (2023).
@@
- Nakamura et al. [2024] K. Nakamura, A. Daido, and Y. Yanase, Orbital effect on the intrinsic superconducting diode effect, Phys. Rev. B 109, 094501 (2024).
@@
- Wang et al. [2024] J. Wang, Y. Jiang, J. J. Wang, and J.-F. Liu, Efficient Josephson diode effect on a two-dimensional topological insulator with asymmetric magnetization, Physical Review B 109, 075412 (2024).
@@
- Cayao et al. [2024] J. Cayao, N. Nagaosa, and Y. Tanaka, Enhancing the Josephson diode effect with Majorana bound states, Physical Review B 109, L081405 (2024).
@@
- Soori [2024] A. Soori, Josephson diode effect in junctions of superconductors with band asymmetric metals, Journal of Physics: Condensed Matter 36, 335303 (2024).
@@
- Volkov et al. [2024] P. A. Volkov, É. Lantagne-Hurtubise, T. Tummuru, S. Plugge, J. H. Pixley, and M. Franz, Josephson diode effects in twisted nodal superconductors, Physical Review B 109, 094518 (2024).
@@
- Cuozzo et al. [2024] J. J. Cuozzo, W. Pan, J. Shabani, and E. Rossi, Microwave-tunable diode effect in asymmetric SQUIDs with topological Josephson junctions, Physical Review Research 6, 023011 (2024).
@@
- Seoane Souto et al. [2024] R. Seoane Souto, M. Leijnse, C. Schrade, M. Valentini, G. Katsaros, and J. Danon, Tuning the Josephson diode response with an ac current, Physical Review Research 6, L022002 (2024).
@@
- Debnath and Dutta [2024] D. Debnath and P. Dutta, Gate-tunable Josephson diode effect in Rashba spin-orbit coupled quantum dot junctions, Physical Review B 109, 174511 (2024).
@@
- Meyer and Houzet [2024] J. S. Meyer and M. Houzet, Josephson diode effect in a ballistic single-channel nanowire (2024), arxiv:2404.01429 [cond-mat] .
@@
- Fracassi et al. [2024] S. Fracassi, S. Traverso, N. T. Ziani, M. Carrega, S. Heun, and M. Sassetti, Anomalous supercurrent and diode effect in locally perturbed topological Josephson junctions (2024), arxiv:2403.17894 [cond-mat] .
@@
- Zazunov et al. [2024] A. Zazunov, J. Rech, T. Jonckheere, B. Grémaud, T. Martin, and R. Egger, Nonreciprocal charge transport and subharmonic structure in voltage-biased Josephson diodes, Physical Review B 109, 024504 (2024).
@@
- Huang et al. [2024] H. Huang, T. de Picoli, and J. I. Väyrynen, Superconducting Diode Effect in Two-dimensional Topological Insulator Edges and Josephson Junctions (2024), arxiv:2404.14566 [cond-mat] .
@@
- Karabassov [2024] T. Karabassov, Anisotropic Josephson Diode Effect in the Topological Hybrid Junctions with the Hexagonal Warping, JETP Letters 119, 316 (2024).
@@
- Correa and Nowak [2024] J. H. Correa and M. P. Nowak, Theory of universal diode effect in three-terminal Josephson junctions (2024), arxiv:2401.16262 [cond-mat] .
@@
- Greco et al. [2024] A. Greco, Q. Pichard, E. Strambini, and F. Giazotto, Double loop dc-SQUID as a tunable Josephson diode (2024), arxiv:2404.05521 [cond-mat] .
@@
- Zhang and Wang [2023] Y. Zhang and Z. Wang, Kramers fulde-ferrell state and superconducting spin diode effect, Phys. Rev. B 107, 224510 (2023).
@@
- Mao et al. [2024] Y. Mao, Q. Yan, Y.-C. Zhuang, and Q.-F. Sun, Universal spin superconducting diode effect from spin-orbit coupling, Phys. Rev. Lett. 132, 216001 (2024).
@@
- Haldane [1988] F. D. M. Haldane, Model for a quantum hall effect without landau levels: Condensed-matter realization of the ”parity anomaly”, Phys. Rev. Lett. 61, 2015 (1988).
@@
- Affleck and Marston [1988] I. Affleck and J. B. Marston, Large-n limit of the heisenberg-hubbard model: Implications for high- superconductors, Phys. Rev. B 37, 3774 (1988).
- Ubbens and Lee [1992] M. U. Ubbens and P. A. Lee, Flux phases in the t-j model, Phys. Rev. B 46, 8434 (1992).
@@
- Lee et al. [2006] P. A. Lee, N. Nagaosa, and X.-G. Wen, Doping a mott insulator: Physics of high-temperature superconductivity, Rev. Mod. Phys. 78, 17 (2006).
@@
- Varma [1997] C. M. Varma, Non-fermi-liquid states and pairing instability of a general model of copper oxide metals, Phys. Rev. B 55, 14554 (1997).
@@
- Varma [1999] C. M. Varma, Pseudogap phase and the quantum-critical point in copper-oxide metals, Phys. Rev. Lett. 83, 3538 (1999).
@@
- Chakravarty et al. [2001] S. Chakravarty, R. B. Laughlin, D. K. Morr, and C. Nayak, Hidden order in the cuprates, Phys. Rev. B 63, 094503 (2001).
@@
- Nayak [2000] C. Nayak, Density-wave states of nonzero angular momentum, Phys. Rev. B 62, 4880 (2000).
@@
- Venderbos [2016] J. W. F. Venderbos, Symmetry analysis of translational symmetry broken density waves: Application to hexagonal lattices in two dimensions, Phys. Rev. B 93, 115107 (2016).
@@
- Liu et al. [2016] J. Liu, S. Y. Park, K. F. Garrity, and D. Vanderbilt, Flux states and topological phases from spontaneous time-reversal symmetry breaking in -based systems, Phys. Rev. Lett. 117, 257201 (2016).
@@
- Classen et al. [2019] L. Classen, C. Honerkamp, and M. M. Scherer, Competing phases of interacting electrons on triangular lattices in moiré heterostructures, Phys. Rev. B 99, 195120 (2019).
@@
- Chen et al. [2024b] Y. Chen, K. Jiang, Y. Zhang, and J. Hu, Flux phases in the extended hubbard model on the triangular lattice, SCIENCE CHINA Physics, Mechanics & Astronomy (2024b).
@@
- Ortiz et al. [2020] B. R. Ortiz, S. M. L. Teicher, Y. Hu, J. L. Zuo, P. M. Sarte, E. C. Schueller, A. M. M. Abeykoon, M. J. Krogstad, S. Rosenkranz, R. Osborn, R. Seshadri, L. Balents, J. He, and S. D. Wilson, : A topological kagome metal with a superconducting ground state, Phys. Rev. Lett. 125, 247002 (2020).
@@
- Ortiz et al. [2021] B. R. Ortiz, S. M. L. Teicher, L. Kautzsch, P. M. Sarte, N. Ratcliff, J. Harter, J. P. C. Ruff, R. Seshadri, and S. D. Wilson, Fermi surface mapping and the nature of charge-density-wave order in the kagome superconductor , Phys. Rev. X 11, 041030 (2021).
@@
- Liang et al. [2021] Z. Liang, X. Hou, F. Zhang, W. Ma, P. Wu, Z. Zhang, F. Yu, J.-J. Ying, K. Jiang, L. Shan, Z. Wang, and X.-H. Chen, Three-dimensional charge density wave and surface-dependent vortex-core states in a kagome superconductor , Phys. Rev. X 11, 031026 (2021).
@@
- Uykur et al. [2021] E. Uykur, B. R. Ortiz, O. Iakutkina, M. Wenzel, S. D. Wilson, M. Dressel, and A. A. Tsirlin, Low-energy optical properties of the nonmagnetic kagome metal , Phys. Rev. B 104, 045130 (2021).
@@
- Shumiya et al. [2021] N. Shumiya, M. S. Hossain, J.-X. Yin, Y.-X. Jiang, B. R. Ortiz, H. Liu, Y. Shi, Q. Yin, H. Lei, S. S. Zhang, G. Chang, Q. Zhang, T. A. Cochran, D. Multer, M. Litskevich, Z.-J. Cheng, X. P. Yang, Z. Guguchia, S. D. Wilson, and M. Z. Hasan, Intrinsic nature of chiral charge order in the kagome superconductor , Phys. Rev. B 104, 035131 (2021).
@@
- Jiang et al. [2021] Y.-X. Jiang, J.-X. Yin, M. M. Denner, N. Shumiya, B. R. Ortiz, G. Xu, Z. Guguchia, J. He, M. S. Hossain, X. Liu, J. Ruff, L. Kautzsch, S. S. Zhang, G. Chang, I. Belopolski, Q. Zhang, T. A. Cochran, D. Multer, M. Litskevich, Z.-J. Cheng, X. P. Yang, Z. Wang, R. Thomale, T. Neupert, S. D. Wilson, and M. Z. Hasan, Unconventional chiral charge order in kagome superconductor KV3Sb5, Nature Materials 20, 1353 (2021).
@@
- Zhao et al. [2021] H. Zhao, H. Li, B. R. Ortiz, S. M. L. Teicher, T. Park, M. Ye, Z. Wang, L. Balents, S. D. Wilson, and I. Zeljkovic, Cascade of correlated electron states in the kagome superconductor CsV3Sb5, Nature 599, 216 (2021).
@@
- Li et al. [2022] H. Li, H. Zhao, B. R. Ortiz, T. Park, M. Ye, L. Balents, Z. Wang, S. D. Wilson, and I. Zeljkovic, Rotation symmetry breaking in the normal state of a kagome superconductor KV3Sb5, Nature Physics 18, 265 (2022).
@@
- Xing et al. [2024] Y. Xing, S. Bae, E. Ritz, F. Yang, T. Birol, A. N. Capa Salinas, B. R. Ortiz, S. D. Wilson, Z. Wang, R. M. Fernandes, and V. Madhavan, Optical manipulation of the charge-density-wave state in rbv3sb5, Nature 631, 60 (2024).
@@
- Feng et al. [2021] X. Feng, K. Jiang, Z. Wang, and J. Hu, Chiral flux phase in the kagome superconductor AV3Sb5, Science Bulletin 66, 1384 (2021).
@@
- Denner et al. [2021] M. M. Denner, R. Thomale, and T. Neupert, Analysis of charge order in the kagome metal (), Phys. Rev. Lett. 127, 217601 (2021).
@@
- Park et al. [2021] T. Park, M. Ye, and L. Balents, Electronic instabilities of kagome metals: Saddle points and Landau theory, Phys. Rev. B 104, 035142 (2021).
@@
- Lin and Nandkishore [2021] Y.-P. Lin and R. M. Nandkishore, Complex charge density waves at van hove singularity on hexagonal lattices: Haldane-model phase diagram and potential realization in the kagome metals AV3Sb5 ( =K, Rb, Cs), Phys. Rev. B 104, 045122 (2021).
@@
- Gu et al. [2022] Y. Gu, Y. Zhang, X. Feng, K. Jiang, and J. Hu, Gapless excitations inside the fully gapped kagome superconductors AV3Sb5, Phys. Rev. B 105, L100502 (2022).
@@
- Dong et al. [2023] J.-W. Dong, Z. Wang, and S. Zhou, Loop-current charge density wave driven by long-range coulomb repulsion on the kagomé lattice, Phys. Rev. B 107, 045127 (2023).
@@
- Zhou and Wang [2022] S. Zhou and Z. Wang, Chern fermi pocket, topological pair density wave, and charge-4e and charge-6e superconductivity in kagomé superconductors, Nature Communications 13, 7288 (2022).
@@
- Tazai et al. [2023] R. Tazai, Y. Yamakawa, and H. Kontani, Charge-loop current order and Z3 nematicity mediated by bond order fluctuations in kagome metals, Nature Communications 14, 7845 (2023).
@@
- Fu et al. [2024] R.-Q. Fu, J. Zhan, M. Dürrnagel, H. Hohmann, R. Thomale, J. Hu, Z. Wang, S. Zhou, and X. Wu, Exotic charge density waves and superconductivity on the Kagome Lattice, arXiv e-prints , arXiv:2405.09451 (2024).
@@
- Colomés and Franz [2018] E. Colomés and M. Franz, Antichiral Edge States in a Modified Haldane Nanoribbon, Phys. Rev. Lett. 120, 086603 (2018).
@@
- Zhu et al. [2013] L. Zhu, V. Aji, and C. M. Varma, Ordered loop current states in bilayer graphene, Physical Review B 87, 035427 (2013).
- Beenakker [1991] C. W. J. Beenakker, Universal limit of critical-current fluctuations in mesoscopic Josephson junctions, Phys. Rev. Lett. 67, 3836 (1991).

### v3

@@ Acknowledgements.
- HuΒ etΒ al. [2023] J.-X.Β Hu, Z.-T.Β Sun, Y.-M.Β Xie,Β andΒ K.Β T.Β Law,Β Josephson diode effect induced by valley polarization in twisted bilayer graphene,Β Phys. Rev. Lett.Β 130,Β 266003 (2023).
@@
- LuΒ etΒ al. [2023] B.Β Lu, S.Β Ikegaya, P.Β Burset, Y.Β Tanaka,Β andΒ N.Β Nagaosa,Β Tunable josephson diode effect on the surface of topological insulators,Β Phys. Rev. Lett.Β 131,Β 096001 (2023).
@@
- AlvaradoΒ etΒ al. [2023] M.Β Alvarado, P.Β Burset,Β andΒ A.Β L.Β Yeyati,Β Intrinsic nonmagnetic josephson junctions in twisted bilayer graphene,Β Phys. Rev. Res.Β 5,Β L032033 (2023).
@@
- NakamuraΒ etΒ al. [2024] K.Β Nakamura, A.Β Daido,Β andΒ Y.Β Yanase,Β Orbital effect on the intrinsic superconducting diode effect,Β Phys. Rev. BΒ 109,Β 094501 (2024).
@@
- WangΒ etΒ al. [2024] J.Β Wang, Y.Β Jiang, J.Β J.Β Wang,Β andΒ J.-F.Β Liu,Β Efficient Josephson diode effect on a two-dimensional topological insulator with asymmetric magnetization,Β Physical Review BΒ 109,Β 075412 (2024).
@@
- CayaoΒ etΒ al. [2024] J.Β Cayao, N.Β Nagaosa,Β andΒ Y.Β Tanaka,Β Enhancing the Josephson diode effect with Majorana bound states,Β Physical Review BΒ 109,Β L081405 (2024).
@@
- Soori [2024] A.Β Soori,Β Josephson diode effect in junctions of superconductors with band asymmetric metals,Β Journal of Physics: Condensed MatterΒ 36,Β 335303 (2024).
@@
- VolkovΒ etΒ al. [2024] P.Β A.Β Volkov, Γ.Β Lantagne-Hurtubise, T.Β Tummuru, S.Β Plugge, J.Β H.Β Pixley,Β andΒ M.Β Franz,Β Josephson diode effects in twisted nodal superconductors,Β Physical Review BΒ 109,Β 094518 (2024).
@@
- CuozzoΒ etΒ al. [2024] J.Β J.Β Cuozzo, W.Β Pan, J.Β Shabani,Β andΒ E.Β Rossi,Β Microwave-tunable diode effect in asymmetric SQUIDs with topological Josephson junctions,Β Physical Review ResearchΒ 6,Β 023011 (2024).
@@
- SeoaneΒ SoutoΒ etΒ al. [2024] R.Β SeoaneΒ Souto, M.Β Leijnse, C.Β Schrade, M.Β Valentini, G.Β Katsaros,Β andΒ J.Β Danon,Β Tuning the Josephson diode response with an ac current,Β Physical Review ResearchΒ 6,Β L022002 (2024).
@@
- DebnathΒ andΒ Dutta [2024] D.Β DebnathΒ andΒ P.Β Dutta,Β Gate-tunable Josephson diode effect in Rashba spin-orbit coupled quantum dot junctions,Β Physical Review BΒ 109,Β 174511 (2024).
@@
- MeyerΒ andΒ Houzet [2024] J.Β S.Β MeyerΒ andΒ M.Β Houzet,Β Josephson diode effect in a ballistic single-channel nanowire (2024),Β arxiv:2404.01429 [cond-mat] .
@@
- FracassiΒ etΒ al. [2024] S.Β Fracassi, S.Β Traverso, N.Β T.Β Ziani, M.Β Carrega, S.Β Heun,Β andΒ M.Β Sassetti,Β Anomalous supercurrent and diode effect in locally perturbed topological Josephson junctions (2024),Β arxiv:2403.17894 [cond-mat] .
@@
- ZazunovΒ etΒ al. [2024] A.Β Zazunov, J.Β Rech, T.Β Jonckheere, B.Β GrΓ©maud, T.Β Martin,Β andΒ R.Β Egger,Β Nonreciprocal charge transport and subharmonic structure in voltage-biased Josephson diodes,Β Physical Review BΒ 109,Β 024504 (2024).
@@
- HuangΒ etΒ al. [2024] H.Β Huang, T.Β de Picoli,Β andΒ J.Β I.Β VΓ€yrynen,Β Superconducting Diode Effect in Two-dimensional Topological Insulator Edges and Josephson Junctions (2024),Β arxiv:2404.14566 [cond-mat] .
@@
- Karabassov [2024] T.Β Karabassov,Β Anisotropic Josephson Diode Effect in the Topological Hybrid Junctions with the Hexagonal Warping,Β JETP LettersΒ 119,Β 316 (2024).
@@
- CorreaΒ andΒ Nowak [2024] J.Β H.Β CorreaΒ andΒ M.Β P.Β Nowak,Β Theory of universal diode effect in three-terminal Josephson junctions (2024),Β arxiv:2401.16262 [cond-mat] .
@@
- GrecoΒ etΒ al. [2024] A.Β Greco, Q.Β Pichard, E.Β Strambini,Β andΒ F.Β Giazotto,Β Double loop dc-SQUID as a tunable Josephson diode (2024),Β arxiv:2404.05521 [cond-mat] .
- MarginedaΒ etΒ al. [2023] D.Β Margineda, A.Β Crippa, E.Β Strambini, Y.Β Fukaya, M.Β T.Β Mercaldo, M.Β Cuoco,Β andΒ F.Β Giazotto,Β Sign reversal diode effect in superconducting dayem nanobridges,Β Communications PhysicsΒ 6,Β 343 (2023).
- MarginedaΒ etΒ al. [2023] D.Β Margineda, A.Β Crippa, E.Β Strambini, L.Β Borgongino, A.Β Paghi, G.Β de Simoni, L.Β Sorba andYuri Fukaya, M.Β T.Β Mercaldo, C.Β Ortix, M.Β Cuoco,Β andΒ F.Β Giazotto,Β Back-action supercurrent diodes,Β arXiv e-printsΒ ,Β arXiv:2311.14503 (2023),Β arXiv:2311.14503 .
- FukayaΒ etΒ al. [2024] Y.Β Fukaya, M.Β T.Β Mercaldo, D.Β Margineda, A.Β Crippa, E.Β Strambini, F.Β Giazotto, C.Β Ortix,Β andΒ M.Β Cuoco,Β Design of supercurrent diode by vortex phase texture,Β arXiv e-printsΒ ,Β arXiv:2403.04421 (2024),Β arXiv:2403.04421 .
- FuΒ etΒ al. [2024] P.-H.Β Fu, Y.Β Xu, S.Β A.Β Yang, C.Β H.Β Lee, Y.Β S.Β Ang,Β andΒ J.-F.Β Liu,Β Field-effect josephson diode via asymmetric spin-momentum locking states,Β Phys. Rev. Appl.Β 21,Β 054057 (2024).
- FuΒ etΒ al. [2024a] P.-H.Β Fu, Y.Β Xu, J.-F.Β Liu, C.Β H.Β Lee,Β andΒ Y.Β Sin Ang,Β Transverse Cooper-Pair Rectifier,Β arXiv e-printsΒ ,Β arXiv:2405.04751 (2024a),Β arXiv:2405.04751 .
@@
- ZhangΒ andΒ Wang [2023] Y.Β ZhangΒ andΒ Z.Β Wang,Β Kramers fulde-ferrell state and superconducting spin diode effect,Β Phys. Rev. BΒ 107,Β 224510 (2023).
@@
- MaoΒ etΒ al. [2024] Y.Β Mao, Q.Β Yan, Y.-C.Β Zhuang,Β andΒ Q.-F.Β Sun,Β Universal spin superconducting diode effect from spin-orbit coupling,Β Phys. Rev. Lett.Β 132,Β 216001 (2024).
@@
- Haldane [1988] F.Β D.Β M.Β Haldane,Β Model for a quantum hall effect without landau levels: Condensed-matter realization of the βparity anomalyβ,Β Phys. Rev. Lett.Β 61,Β 2015 (1988).
@@
- AffleckΒ andΒ Marston [1988] I.Β AffleckΒ andΒ J.Β B.Β Marston,Β Large-n limit of the heisenberg-hubbard model: Implications for high- superconductors,Β Phys. Rev. BΒ 37,Β 3774 (1988).
- UbbensΒ andΒ Lee [1992] M.Β U.Β UbbensΒ andΒ P.Β A.Β Lee,Β Flux phases in the t-j model,Β Phys. Rev. BΒ 46,Β 8434 (1992).
@@
- LeeΒ etΒ al. [2006] P.Β A.Β Lee, N.Β Nagaosa,Β andΒ X.-G.Β Wen,Β Doping a mott insulator: Physics of high-temperature superconductivity,Β Rev. Mod. Phys.Β 78,Β 17 (2006).
@@
- Varma [1997] C.Β M.Β Varma,Β Non-fermi-liquid states and pairing instability of a general model of copper oxide metals,Β Phys. Rev. BΒ 55,Β 14554 (1997).
@@
- Varma [1999] C.Β M.Β Varma,Β Pseudogap phase and the quantum-critical point in copper-oxide metals,Β Phys. Rev. Lett.Β 83,Β 3538 (1999).
@@
- ChakravartyΒ etΒ al. [2001] S.Β Chakravarty, R.Β B.Β Laughlin, D.Β K.Β Morr,Β andΒ C.Β Nayak,Β Hidden order in the cuprates,Β Phys. Rev. BΒ 63,Β 094503 (2001).
@@
- Nayak [2000] C.Β Nayak,Β Density-wave states of nonzero angular momentum,Β Phys. Rev. BΒ 62,Β 4880 (2000).
@@
- Venderbos [2016] J.Β W.Β F.Β Venderbos,Β Symmetry analysis of translational symmetry broken density waves: Application to hexagonal lattices in two dimensions,Β Phys. Rev. BΒ 93,Β 115107 (2016).
@@
- LiuΒ etΒ al. [2016] J.Β Liu, S.Β Y.Β Park, K.Β F.Β Garrity,Β andΒ D.Β Vanderbilt,Β Flux states and topological phases from spontaneous time-reversal symmetry breaking in -based systems,Β Phys. Rev. Lett.Β 117,Β 257201 (2016).
@@
- ClassenΒ etΒ al. [2019] L.Β Classen, C.Β Honerkamp,Β andΒ M.Β M.Β Scherer,Β Competing phases of interacting electrons on triangular lattices in moirΓ© heterostructures,Β Phys. Rev. BΒ 99,Β 195120 (2019).
@@
- ChenΒ etΒ al. [2024b] Y.Β Chen, K.Β Jiang, Y.Β Zhang,Β andΒ J.Β Hu,Β Flux phases in the extended hubbard model on the triangular lattice,Β Science China Physics, Mechanics & AstronomyΒ 67,Β 297211 (2024b).
@@
- OrtizΒ etΒ al. [2020] B.Β R.Β Ortiz, S.Β M.Β L.Β Teicher, Y.Β Hu, J.Β L.Β Zuo, P.Β M.Β Sarte, E.Β C.Β Schueller, A.Β M.Β M.Β Abeykoon, M.Β J.Β Krogstad, S.Β Rosenkranz, R.Β Osborn, R.Β Seshadri, L.Β Balents, J.Β He,Β andΒ S.Β D.Β Wilson,Β : A topological kagome metal with a superconducting ground state,Β Phys. Rev. Lett.Β 125,Β 247002 (2020).
@@
- OrtizΒ etΒ al. [2021] B.Β R.Β Ortiz, S.Β M.Β L.Β Teicher, L.Β Kautzsch, P.Β M.Β Sarte, N.Β Ratcliff, J.Β Harter, J.Β P.Β C.Β Ruff, R.Β Seshadri,Β andΒ S.Β D.Β Wilson,Β Fermi surface mapping and the nature of charge-density-wave order in the kagome superconductor ,Β Phys. Rev. XΒ 11,Β 041030 (2021).
@@
- LiangΒ etΒ al. [2021] Z.Β Liang, X.Β Hou, F.Β Zhang, W.Β Ma, P.Β Wu, Z.Β Zhang, F.Β Yu, J.-J.Β Ying, K.Β Jiang, L.Β Shan, Z.Β Wang,Β andΒ X.-H.Β Chen,Β Three-dimensional charge density wave and surface-dependent vortex-core states in a kagome superconductor ,Β Phys. Rev. XΒ 11,Β 031026 (2021).
@@
- UykurΒ etΒ al. [2021] E.Β Uykur, B.Β R.Β Ortiz, O.Β Iakutkina, M.Β Wenzel, S.Β D.Β Wilson, M.Β Dressel,Β andΒ A.Β A.Β Tsirlin,Β Low-energy optical properties of the nonmagnetic kagome metal ,Β Phys. Rev. BΒ 104,Β 045130 (2021).
@@
- ShumiyaΒ etΒ al. [2021] N.Β Shumiya, M.Β S.Β Hossain, J.-X.Β Yin, Y.-X.Β Jiang, B.Β R.Β Ortiz, H.Β Liu, Y.Β Shi, Q.Β Yin, H.Β Lei, S.Β S.Β Zhang, G.Β Chang, Q.Β Zhang, T.Β A.Β Cochran, D.Β Multer, M.Β Litskevich, Z.-J.Β Cheng, X.Β P.Β Yang, Z.Β Guguchia, S.Β D.Β Wilson,Β andΒ M.Β Z.Β Hasan,Β Intrinsic nature of chiral charge order in the kagome superconductor ,Β Phys. Rev. BΒ 104,Β 035131 (2021).
@@
- JiangΒ etΒ al. [2021] Y.-X.Β Jiang, J.-X.Β Yin, M.Β M.Β Denner, N.Β Shumiya, B.Β R.Β Ortiz, G.Β Xu, Z.Β Guguchia, J.Β He, M.Β S.Β Hossain, X.Β Liu, J.Β Ruff, L.Β Kautzsch, S.Β S.Β Zhang, G.Β Chang, I.Β Belopolski, Q.Β Zhang, T.Β A.Β Cochran, D.Β Multer, M.Β Litskevich, Z.-J.Β Cheng, X.Β P.Β Yang, Z.Β Wang, R.Β Thomale, T.Β Neupert, S.Β D.Β Wilson,Β andΒ M.Β Z.Β Hasan,Β Unconventional chiral charge order in kagome superconductor KV3Sb5,Β Nature MaterialsΒ 20,Β 1353 (2021).
@@
- ZhaoΒ etΒ al. [2021] H.Β Zhao, H.Β Li, B.Β R.Β Ortiz, S.Β M.Β L.Β Teicher, T.Β Park, M.Β Ye, Z.Β Wang, L.Β Balents, S.Β D.Β Wilson,Β andΒ I.Β Zeljkovic,Β Cascade of correlated electron states in the kagome superconductor CsV3Sb5,Β NatureΒ 599,Β 216 (2021).
@@
- LiΒ etΒ al. [2022] H.Β Li, H.Β Zhao, B.Β R.Β Ortiz, T.Β Park, M.Β Ye, L.Β Balents, Z.Β Wang, S.Β D.Β Wilson,Β andΒ I.Β Zeljkovic,Β Rotation symmetry breaking in the normal state of a kagome superconductor KV3Sb5,Β Nature PhysicsΒ 18,Β 265 (2022).
@@
- XingΒ etΒ al. [2024] Y.Β Xing, S.Β Bae, E.Β Ritz, F.Β Yang, T.Β Birol, A.Β N.Β CapaΒ Salinas, B.Β R.Β Ortiz, S.Β D.Β Wilson, Z.Β Wang, R.Β M.Β Fernandes,Β andΒ V.Β Madhavan,Β Optical manipulation of the charge-density-wave state in rbv3sb5,Β NatureΒ 631,Β 60 (2024).
@@
- FengΒ etΒ al. [2021] X.Β Feng, K.Β Jiang, Z.Β Wang,Β andΒ J.Β Hu,Β Chiral flux phase in the kagome superconductor AV3Sb5,Β Science BulletinΒ 66,Β 1384 (2021).
@@
- DennerΒ etΒ al. [2021] M.Β M.Β Denner, R.Β Thomale,Β andΒ T.Β Neupert,Β Analysis of charge order in the kagome metal (),Β Phys. Rev. Lett.Β 127,Β 217601 (2021).
@@
- ParkΒ etΒ al. [2021] T.Β Park, M.Β Ye,Β andΒ L.Β Balents,Β Electronic instabilities of kagome metals: Saddle points and Landau theory,Β Phys. Rev. BΒ 104,Β 035142 (2021).
@@
- LinΒ andΒ Nandkishore [2021] Y.-P.Β LinΒ andΒ R.Β M.Β Nandkishore,Β Complex charge density waves at van hove singularity on hexagonal lattices: Haldane-model phase diagram and potential realization in the kagome metals AV3Sb5 ( =K, Rb, Cs),Β Phys. Rev. BΒ 104,Β 045122 (2021).
@@
- GuΒ etΒ al. [2022] Y.Β Gu, Y.Β Zhang, X.Β Feng, K.Β Jiang,Β andΒ J.Β Hu,Β Gapless excitations inside the fully gapped kagome superconductors AV3Sb5,Β Phys. Rev. BΒ 105,Β L100502 (2022).
@@
- DongΒ etΒ al. [2023] J.-W.Β Dong, Z.Β Wang,Β andΒ S.Β Zhou,Β Loop-current charge density wave driven by long-range coulomb repulsion on the kagomΓ© lattice,Β Phys. Rev. BΒ 107,Β 045127 (2023).
@@
- ZhouΒ andΒ Wang [2022] S.Β ZhouΒ andΒ Z.Β Wang,Β Chern fermi pocket, topological pair density wave, and charge-4e and charge-6e superconductivity in kagomΓ© superconductors,Β Nature CommunicationsΒ 13,Β 7288 (2022).
@@
- TazaiΒ etΒ al. [2023] R.Β Tazai, Y.Β Yamakawa,Β andΒ H.Β Kontani,Β Charge-loop current order and Z3 nematicity mediated by bond order fluctuations in kagome metals,Β Nature CommunicationsΒ 14,Β 7845 (2023).
@@
- FuΒ etΒ al. [2024b] R.-Q.Β Fu, J.Β Zhan, M.Β DΓΌrrnagel, H.Β Hohmann, R.Β Thomale, J.Β Hu, Z.Β Wang, S.Β Zhou,Β andΒ X.Β Wu,Β Exotic charge density waves and superconductivity on the Kagome Lattice,Β arXiv e-printsΒ ,Β arXiv:2405.09451 (2024b).
@@
- ColomΓ©sΒ andΒ Franz [2018] E.Β ColomΓ©sΒ andΒ M.Β Franz,Β Antichiral Edge States in a Modified Haldane Nanoribbon,Β Phys. Rev. Lett.Β 120,Β 086603 (2018).
@@
- ZhuΒ etΒ al. [2013] L.Β Zhu, V.Β Aji,Β andΒ C.Β M.Β Varma,Β Ordered loop current states in bilayer graphene,Β Physical Review BΒ 87,Β 035427 (2013).
@@
- Beenakker [1991] C.Β W.Β J.Β Beenakker,Β Universal limit of critical-current fluctuations in mesoscopic Josephson junctions,Β Phys. Rev. Lett.Β 67,Β 3836 (1991).